How do you solve (x-1/2)²=1 i've tried to square root both sides i put them under the same denominator which is 2. Then i dropped them and i got 2x-1=(square root)1 +1. Then i added +1 to each side and my final answer was x=1+(square root)1 (all divided by) 2 BUT it is not one of the answers on my homework sheet. Hope you can help, Thank You!
Answer by stanbon(75887)

(x-1/2)^2 = 1
------
[(x-(1/2))^2-1 = 0
---
Factor:
[(x-(1/2))-1][(x-(1/2))+1] = -
----
(x-(1/2) = 1 or (x-(1/2) = -1
-----
x = 3/2 or x = -1/2
========================
Cheers,
Stan H.